OpenStudy (anonymous):

Help ASAP! Trying to get exempt from the final. 2. Jupiter has a radius of about 11.0 earth radii, and a mass 318 times the earth’s mass. Estimate the gravitational field on the surface of Jupiter compared to that on the earth. Show your work.

OpenStudy (anonymous):

Acceleration due to gravity is \[\frac{GM}{r^2}\] you want to work out the ratio of that of Jupiter to Earth \[\frac{GM_J}{r_J^2} \div \frac{GM_E}{r_E^2}\] in the question you're told \[r_J = 11\times r_E\] \[M_J = 318 \times r_E\] I'm sure you're clever enough to plug in the numbers and cancel out the appropriate variables.

OpenStudy (anonymous):

Whoops, I meant \[M_J = 318\times M_E\]

OpenStudy (anonymous):

got it thanks (: